OpenStudy (anonymous):

a^2 +ac +ba +bc over a-b ÷ a-b

OpenStudy (anonymous):

does it help to know that \((a-b)\div (a-b)=1\)?

OpenStudy (anonymous):

It is supposed to say a^2 +ac +ba +bc over a-b ÷ a-c. My fault

OpenStudy (mathstudent55):

Is this the problem? \(\large \dfrac{a^2 + ac + ba + bc}{\frac{a-b}{a-c} }\)

OpenStudy (mathstudent55):

If so, remember that dividing by a fraction is the same as multiplying by its reciprocal. Also, since the numerator has four terms, try factoring it by grouping.

OpenStudy (anonymous):

No, but how did you do the big fraction bar?

OpenStudy (mathstudent55):

\dfrac is a big fraction bar \frac is a small fraction bar

OpenStudy (mathstudent55):

Sorry, gtg.